The sum of the two digit number is 8, and their difference is 6. If the value of the number is less than 50, what is the number?
Could someone help me please?

Answer by stanbon(75887)

The sum of the two digit number is 8, and their difference is 6. If the value of the number is less than 50, what is the number?
--------------------
Let the two-digit number be 10t+u where t is the tens and u is the units digit.
-----------------------
EQUATION:
u + t = 8
u - t = 6
------------
Add to solve for "u":
2u = 14
u =7
Substitute in u+t=8 to get t=1
----------------
The number is 10*1+7 = 17
==============================
Cheers,
Stan H.